# Security Weaknesses of Copilot Generated Code in GitHub

### v2

@@ 6. Threats to Validity
Construct Validity: This study has three threats to construct validity: (1) Using keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub. The results obtained through the keyword-based search may not cover all the code snippets generated by Copilot on GitHub. We tried to mitigate this threat by constantly and iteratively refining the keywords and using synonyms. (2) Manual data filtering – We manually screened the results obtained from the keyword-based search by analyzing the comments, tags, and other metadata of the code snippets to determine whether they were generated by Copilot. Since this process was manually done, it may have been influenced by personal bias. In this regard, two authors conducted the pilot experiment independently to minimize the impact on the construct validity. (3) Manual association of CWEs – We manually associated the warning messages reported by the static analysis tools with a particular CWE, which may introduce personal subjective bias, threatening the construct validity. We employed two measures to mitigate this threat. First, since the list of CWEs is a tree structure with interconnections between them, we first matched the warning messages to a higher-level CWE, and further checked whether we can match the warning messages to a lower-level CWE with more specific definition. Second, to mitigate the personal bias, two authors independently assigned each security issue description a CWE ID. In case of disagreement, the two authors discussed it with the assessment by a third author (a security expert).
@@
External Validity: Our dataset consists of Copilot-generated code snippets collected from open-source projects on GitHub. During the filtering process, we excluded code that utilized Copilot to solve algorithmic problems, aiming to ensure that the collected data genuinely reflected real-world production environments. Since the data from GitHub are not diversified enough, we had a higher number of code snippets originating from the Game projects. This could result in a lack of comprehensiveness in the security scenarios involved. The peculiarity of the data source may make the dataset incomplete, thereby threatening the external validity of the results. Furthermore, we acknowledge the need to collect more diverse code snippets from different platforms to increase the generalizability of the results. We will consider adopting more diversified ways or platforms to collect code. Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false positives in the scan results (as the case with static analysis, in general, (Sui et al., 2020; Kang et al., 2022)). Although we used two widely used static analysis tools to increase the comprehensiveness of the scans and manually checked the results of the tool scans, the results may suffer from incompleteness.
@@
Reliability: We used multiple automated static analysis tools to analyze the Copilot-generated code snippets to improve security weaknesses detection. Developers have widely used these automated tools. The querying mechanism of these tools ensures that the scan results remain consistent when used multiple times. In addition, we performed two rounds of scanning with two tools for security checks on each code snippet, intending to complement the results of one tool with the other. By implementing these measures, we believe that our research results are reliable and these threats to reliability are mitigated.
@@ 7. Conclusions
Automatic code generation and recommendation have been an active research area due to the advancement of AI and, specifically, LLMs. AI code generation tools, such as Copilot, can significantly improve developers’ development efficiency, but can also introduce vulnerabilities and security risks. In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot-generated code found in public GitHub projects. We identified 452 code snippets generated by Copilot from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools. This study aims to help developers understand the security risks of weaknesses introduced in the code generated by Copilot (and potentially similar code generation tools). Our results show that (1) 29.8% of the 452 Copilot-generated code snippets contain security weaknesses. Developers have a high risk of raising security issues when using Copilot, regardless of the programming language, so security checks are necessary. (2) The detected security weaknesses are diverse in nature and are associated with 38 different CWEs. Developers face a variety of development scenarios and application environments in production and need the appropriate security awareness and skills. (3) Among these CWEs, eight appear in the MITRE CWE Top-25 list, and six belong to the Stubborn Weaknesses, demonstrating their severity.
@@
Data availability
We have shared the link to our dataset in the reference (Fu et al., 2024).

### v3

@@ 6. Threats to Validity
Construct Validity: This study has three threats to construct validity: (1) Using keyword-based search – We used a keyword-based search to collect relevant code snippets from GitHub. The results obtained through the keyword-based search may not cover all the code snippets generated by Copilot and other tools on GitHub. We tried to mitigate this threat by constantly and iteratively refining the keywords and using synonyms. (2) Manual data filtering – We manually screened the results obtained from the keyword-based search by analyzing the comments, tags, and other metadata of the code snippets to determine whether they were generated by Copilot or other tools. Since this process was manually done, it may have been influenced by personal bias. In this regard, two authors conducted the experiment independently to minimize the impact on the construct validity. (3) Manual association of CWEs – We manually associated the warning messages reported by the static analysis tools with a particular CWE, which may introduce personal subjective bias, threatening the construct validity. We employed two measures to mitigate this threat. First, since the list of CWEs is a tree structure with interconnections between them, we first matched the warning messages to a higher-level CWE and further checked whether we can match the warning messages to a lower-level CWE with a more specific definition. We refer to the description of the test suite and CWE coverage in the static analysis tool documentation. Second, to mitigate the personal bias, two authors independently assigned each security weakness description a CWE ID within a period of ten days. In case of disagreement, the two authors discussed it with the assessment by a third author (a security expert).
@@
External Validity: Our dataset consists of the code snippets generated by Copilot and other code generation tools collected from open-source projects on GitHub. During the filtering process, we excluded code that utilized AI code generation tools to solve simple programming practice problems, aiming to ensure that the collected data genuinely reflected open source development on GitHub. Since the data from GitHub are not diversified enough, we had a higher number of code snippets originating from the Game projects. This could result in a lack of comprehensiveness in the security scenarios involved. The peculiarity of the data source may make the dataset incomplete, thereby threatening the external validity of the results. Furthermore, we acknowledge the need to collect more diverse code snippets from different platforms to increase the generalizability of the results. We will consider adopting more diversified ways or platforms to collect code.Additionally, due to the limitations of static analysis tools themselves, these tools could not scan all CWEs, and there is a degree of false positives in the analysis results (as the case with static analysis, in general, (Sui et al., 2020; Kang et al., 2022)). Although we used two widely used static analysis tools to increase the comprehensiveness of the scans and manually checked the results of the tool scans, the results may suffer from incompleteness.
@@
Reliability: We used multiple automated static analysis tools to analyze the Copilot-generated code snippets to improve security weaknesses detection. Developers have widely used these automated tools. The querying mechanism of these tools ensures that the analysis results remain consistent when used multiple times. In addition, we performed two rounds of analyzing with two tools for security checks on each code snippet, intending to complement the results of one tool with the other. By implementing these measures, we believe that our research results are reliable and these threats to reliability are mitigated.
@@ 7. Conclusions
Automatic code generation and recommendation have been an active research area due to the advancement of AI and, specifically, LLMs. AI code generation tools, such as Copilot, can significantly improve developers’ development efficiency but can also introduce vulnerabilities and security risks. Existing studies on the security of AI-generated code mainly focused on the security issues in the generated code using crafted scenarios, and potential security weaknesses of AI-generated code in practical scenarios in open source development environment (e.g., GitHub) have not been fully considered. In this paper, we present the results of an empirical study to analyze security weaknesses in Copilot-generated code found in public GitHub projects. We identified 733 code snippets generated by Copilot and other tools (i.e., CodeWhisperer and Codeium) from GitHub projects and analyzed those snippets for security weaknesses using static analysis tools. We also examined using Copilot Chat to fix security issues in Copilot-generated code. This study aims to help developers understand the security risks of weaknesses introduced in the code generated by Copilot (and potentially similar code generation tools). Our results show that (1) around 30% of the 733 generated code snippets contain security weaknesses. Developers have a high risk of raising security weaknesses when using Copilot or other AI code generation tools, regardless of the programming language, and therefore, applying appropriate security checks is necessary. (2) The detected security weaknesses are diverse in nature and are associated with 43 different CWEs. Developers encounter a variety of development scenarios and production environments, requiring appropriate security awareness and skills. (3) Among these CWEs, eight appear in the 2023 CWE Top-25 list, and six belong to the Stubborn Weaknesses, indicating high severity. (4) Copilot Chat can help fix many of the security issues in Copilot-generated code, but its success rate varies across different CWEs. (5) Providing Copilot Chat with a warning message from the static analysis tool leads to a better fix.